An arithmetic sequence whose fourth entry is 2 and whose seventh entry is 7
Answer by richwmiller(17219)

An = A1 + (n - 1)d
7= A1 + 6d
2= A1 + 3d
5=3d
5/3=d
2=A1+3*5/3
2=A1+5
A1=-3
An = A1 + (n - 1)d
An = -3 + 5/3(n - 1)
An = -3 + 5/3n - 5/3
An = (5n - 14)/3